# Block: Balancing Load in LLM Serving with Context, Knowledge and Predictive Scheduling

### v1

@@ Abstract.
This paper presents Block, a distributed scheduling framework designed to optimize load balancing and auto-provisioning across instances in large language model serving frameworks by leveraging contextual information from incoming requests. Unlike popular model serving systems that rely on monolithic and heuristic task schedulers, Block operates as a fully distributed, stateless, and predictive scheduling system to achieve low overhead, reliability, and scalability. It leverages the deterministic and predictable characteristics of LLM inferences, such as host configurations, response lengths, and hardware performance, to make scheduling decisions based on accurately predicted metrics. Evaluation on a 12 GPUs cluster shows that Block significantly outperforms heuristic schedulers, boosting serving capacity by up to 16.7% and reducing P99 tail latency by up to 49.5%. These performance gains remain consistent across diverse models, workloads and configurations. Code and data are open-sourced.
@@ 1. Introduction
The rise of Large Language Models (LLMs) like GPT4 (OpenAI et al., 2024), Llama  (Touvron et al., 2023), and Gemini (Team et al., 2024) have revolutionized modern applications, such as chatbots (OpenAI et al., 2024), virtual assistants  (Guan et al., 2023), code generation  (Jiang et al., 2024), and creative writing  (Gómez-Rodríguez and Williams, 2023). This places immense pressure on LLM inference serving systems, which must meet stringent latency requirements for a seamless user experience and maximize throughput to handle growing demand (Shen and Sen, 2025; Cheng et al., 2025; Kwon et al., 2023; Yu et al., 2022). To this end, in recent years, we observe significant progress in the development of LLM serving systems. Several key techniques, such as continuous batching (Yu et al., 2022), page attention (Kwon et al., 2023), chunked prefill (Agrawal et al., 2023), and Flash attention (Dao, 2023), have been developed, demonstrating significantly improvements in both latency and throughput.
@@
However, current LLM inference process is often characterized as unpredictable (Sun et al., 2024). For example, the autoregressive nature of LLMs, which generates tokens sequentially based on preceding ones until a stop signal is reached, leads to variable response lengths and decoding steps (Schuurmans et al., 2024). Page attention (Kwon et al., 2023), which dynamically allocates memory resources and allows for request preemption, further contributes to the dynamic nature of runtime memory consumption. Further, latency of each decoding step also exhibits high variance, due to dynamic batch size.
@@
The above uncertainties challenge both scheduling LLM requests and scaling model instances, since widely used runtime metrics such as latency, throughput and memory, no longer accurately present end-to-end execution loads in the case of LLM serving. Further, scheduled requests with unexpectedly long responses can further increase memory load on overloaded hosts and block subsequent requests from being launched (Sun et al., 2024). Most current multi-instance serving frameworks on production, such as (nvi, [n. d.]; Aminabadi et al., 2022), employ scheduling heuristics like round-robin for request dispatching which offers no guarantee on scheduling performance. Alternative solutions, such as Llumnix (Sun et al., 2024), schedule tasks in alignment with dynamic rebalancing through live migration, demonstrating improvements in overall cluster serving performance. However, such solutions require transferring requests’ KV cache (Kang et al., 2024) across instances with extra cost of network bandwidth and memory, which may not be preferable in a resource-limited cluster with high serving pressure.
@@
Recent studies indicate that the uncertainty of inference behavior can be mitigated with the help of purposely built assistant models. For example, the length of responses and duration of execution can be accurately predicted with a pretrained regression model (Zheng et al., 2023b) and via the inference simulation framework (Agrawal et al., 2024a). This presents an unexplored opportunity to improve load balancing in inference systems by exploiting such approaches. For instance, consider the request to “explain the theory of relativity”, which involves short prompts but generates lengthy responses. Predicting a request’s length allows the scheduler to proactively route it to less-loaded devices, which would improve cluster balance.
@@
In this paper, we present Block, a novel decentralized scheduler for LLM inference clusters. The key novelty of Block relies on leveraging queries’ context and utilizes static properties of LLM serving backends, such as hardware capability and serving framework batching strategy, to predicatively schedule inference requests to serving instances.
@@
Block operates in the following way. A lightweight LLM-based regression model is first applied to estimate the length of response based on the request contexts. Then, a simulation framework is applied to predict key target metrics for each request. Finally, the Block task scheduler dispatches requests based on these simulation-based predictions. Following the above steps, Block could estimate real serving load and metrics for each request during scheduling, thereby improving overall performance by maximizing both throughput and resource utilization while minimizing both latency and preemptions. Additionally, our work identifies that such simulation-based prediction provides an efficient approach for auto-provisioning.
@@
We evaluate Block against widely-used heuristic dispatchers on a 12 GPUs cluster with real-world datasets and LLMs. Results demonstrate Block’s improved performance: it increases serving capacity by up to 16.7% and boosts throughput by up to 4.4% against the Llumnix dispatcher. Critically, Block reduces average request latency by 19.9-45.8% and P99 tail latency by 12.6-49.5% compared to baselines. Block’s performance gains are more pronounced when looking at the Time-To-First-Token (TTFT) where average and P99 TTFT are reduced by 88.1-97.0% and 78.6-94.5% respectively. Further, Block improves resource utilization and achieves P99 20.1% latency reduction when using prediction for auto-provisioning.
@@
In summary, this paper makes the following contributions. First, we thoroughly motivate our approach to using predictive scheduling for LLM serving with the integration of response length estimation and inference simulation techniques in §2. Second, we design (§4) and implement, (§5) Block, a distributed scheduling framework with predictive scheduling. Third, we conduct a comprehensive evaluation on a 12 GPUs cluster with real-world datasets/models to demonstrate the superior effectiveness of Block on load balancing and auto-provisioning, as presented in §6. Finally, Block paves the way for a new class of predictive LLM schedulers for online serving. Block’s code and dataset are open-sourced at https://github.com/AKafakA/Block.
@@ 2. Background
Modern large language models (LLMs) are mainly built on transformer architectures (Vaswani et al., 2017) to process input sequences. LLM queries typically consist of strings of variable length known as prompts or contexts. Queries are first tokenized and converted into a sequence of embeddings, which are then processed by transformer blocks (Vaswani et al., 2017) by applying attention mechanisms in conjunction with multilayer perceptrons to project them into a specified space. The procedure of processing prompts is known as encoding.
@@
Due to autoregressive generation, the decoding phase cannot be accelerated through parallelization and requires to cache all tokens until completed. The only exception occurs at the first decoded token, as it only uses the prompt as input without depending on a prior stage. This phase, which involves computing the KV cache for the context sequence, is referred to as the prefill phase, in contrast to the subsequent sequential decoding steps known as the decode phase. To this end, prefill is considered computed-bound while decoding is regarded as memory-bandwidth-bound (Kamath et al., 2024).
Continuous Batching and Page Attention
@@
Figure 1. Page Attention and Preemption
A key challenge in continuous batching is memory allocation. Since response length is unknown at runtime, memory must be pre-allocated for the maximum possible sequence length and leads to memory fragmentation and waste, especially for requests generating shorter responses. This inefficiency on memory utilization necessitates a more dynamic memory management solution. To address this challenge, page attention (Kwon et al., 2023) is proposed and open-sourced through the vLLM serving framework. In the case of page attention, instead of reserving maximum memory or risking an OutOfMemory error, vLLM divides GPU memory into fixed-size memory blocks and maintains a page table to track the locations of noncontinuous physical memory blocks for KV cache associated with each request. As illustrated in Figure 1, preemption occurs when memory is insufficient for the next decoding step. The newest request in the batch is preempted, returned to the head of the waiting queue, and its memory blocks are released for the remaining requests. Once memory becomes available, typically from completed requests, the preempted request resumes, requiring its KV cache to be recomputed. This dynamic memory allocation significantly improves the LLM serving instance’s capacity and throughput, making vLLM a leading framework in research and industry applications.
@@ 3. Related Work and Motivation
This inherent unpredictability in resource requirements and execution behavior makes dispatching requests across LLM serving instances more challenging than general dynamic task scheduling (Savvas and Kechadi, 2004). For example, a standard scheduler like Kubernetes might filter compute nodes based on initially available resources and then selects the node with the lowest current utilization. However, such static filtering and simple utilization metrics are often inadequate for LLM inference workloads because memory demands of requests are dynamic, varying with the decoding length. Additionally, the actual performance when executing a request can still be impacted by interference from other colocated requests (Sun et al., 2024). Furthermore, unexpected preemption can also occur leading to serving request reset and successors blocked.
@@
To ensure a stable application and a good user experience in serving clusters with variable workloads, effective resource management must go beyond scheduling to include auto-provision to handle peak loads and guarantees performance under established SLOs. Although auto-provision is highly desirable and comes with well-known challenges often faced in other areas with similar scheduling characteristics, yet it is often overlooked in LLM serving systems design. Asynchronous cold start prevalent in serverless computing (Liu et al., 2023; Joosen et al., 2024) is such an example. After new instances are provisioned, scheduled functions continue to be executed on heavily loaded instances, resulting in prolonged latencies and maintain high instance load, although the newly provisioned instances are initially under-utilized. Such load imbalance could causing resource wastage and high tail latencies. Similar asynchronous cold start also manifests in LLM serving clusters as in the case of serverless computing systems, they both employ a shared scheduling paradigm, wherein each instance maintains a queue of pending tasks, and auto-provisioning decisions are based on runtime-monitored metrics.
Llumnix (Sun et al., 2024) mitigates such imbalances through dynamic re-balancing. Initially, it dispatches requests using a heuristic scheduler. Subsequently, Llumnix performs continuous dynamic load rebalancing across instances by migrating active requests along with their KV caches. While dynamic rebalancing can mitigate runtime load imbalances, it requires significant GPU memory and inter-GPU network bandwidth to transfer associated token caches, impacting model parallelism performance. This resource contention can even intensify either: 1) when auto-provisioning is triggered due to cold start issues, to re-balance load between heavily utilized existing hosts and newly provisioned ones; or 2) when other network-heavy features are applied, such as P-D disaggregation and tensor/pipeline parallelism (Shoeybi et al., 2020; Huang et al., 2019).
@@
With the rapid development and expansion of new LLM models, devices, and use cases, setting up an appropriate LLM serving cluster has become a critical yet challenging task for developers, particularly due to the high trial-and-error costs associated with GPU pricing. To address this, Vidur, the first LLM cluster simulation framework, has been proposed (Agrawal et al., 2024a). It aims to reduce potential hardware costs when searching for the optimal cluster configuration based on a given model, trace, and Service Level Objective (SLO) requirements. The insight behind Vidur is that since the local scheduler and batching logic are deterministic, and response lengths are typically available in the replay trace used to evaluate cluster performance, it is feasible to simulate the entire replay process if the execution time for each batch can be inferred. Additionally, by profiling low-level operators like attention and linear projection for a specific GPU, Vidur trains linear models to interpolate execution times for various batches. It finally achieves less than 9% error for key metrics as throughput and latencies across a diverse set of models and GPU types. Besides, SimAI (Wang et al., 2025) also provides modeling and simulation for the entire LLM training process with average 98% accuracy.
@@
While some methods predict inference performance assuming known response lengths, alternative approaches also predict the initially unknown lengths to enhance inference itself. Recognizing that the variable length of responses is a key challenge, these techniques leverage prediction to improve local scheduler efficiency. For instance, Sequence Scheduling (Zheng et al., 2023b) proposes instruction-tuning a smaller LLM specifically to predict the response length of the main serving LLM. This prediction allows their local scheduler to be optimized by grouping requests with similar expected lengths into the same batch. Such grouping minimizes computational waste and improves single instances’ throughput by 85%.
Similarly, LightLLM (Gong et al., 2025) predicts request output lengths using historical distributions. Based on these predictions, it infers the peak memory requirement for running batches, enabling the local scheduler to proactively avoid preemption. In the context of global scheduling, length estimation has also been utilized with recent works, primarily as a filtering mechanism prior to scheduling tasks. For instance, TetriServe (Hu et al., 2024) deploys a lightweight length-estimation model to filter out decoding candidate instances without sufficient GPU memory to accommodate the estimated number of tokens and then applies the Power-of-Two method (Richa et al., 2000) to select decoding instances with fewer pending requests. DynamoLLM (Stojkovic et al., 2024) classifies requests into three distinct pools based on their estimated lengths, scheduling them separately with associated host pools. To summarize, our insight is that by combining simulation techniques with length prediction, the unpredictability can be largely eliminated, which in turn creates an opportunity to improve multi-instance LLM serving performance.
@@ 4. Block System Design
Block, as shown in Figure 4, comprises of four services: query length tagger, global scheduler, predictor, and inference framework backend, to handle requests distribution and results collection. The query length tagger service is the entry point of the system and is designed to predict and label the anticipated response lengths to requests. LLM Requests are executed on the GPUs running under an inference framework such as (Kwon et al., 2023; Zheng et al., 2024; Gong et al., 2025). The global scheduler is tasked with making scheduling and auto-provisioning decisions and directing requests to appropriate model instances to balance the load. When a new query is received, the global scheduler first disseminates prediction requests to the predictor services in order to collect instance statuses/predictions for scheduling. Then, it dispatches requests to the selected model instances. Further, the predictor service performs as a sidecar to predict the target metrics. The Block framework is designed to be agnostic to models, hardware, inference frameworks, and scheduling strategies.

### v2

@@ Abstract.
This paper presents Block, a distributed scheduling framework designed to optimize load balancing and auto-provisioning across instances in large language model serving frameworks by leveraging contextual information from incoming requests. Unlike popular model serving systems that rely on monolithic and heuristic task schedulers, Block operates as a fully distributed, stateless, and predictive scheduling system to achieve low overhead, reliability, and scalability. It leverages the deterministic and predictable characteristics of LLM inference, such as host configurations, response lengths, and hardware performance, to make scheduling decisions based on accurately predicted metrics. Evaluation on a 12-GPU cluster shows that Block significantly outperforms heuristic schedulers, boosting serving capacity by up to 16.7% and reducing P99 tail latency by up to 49.5%. These performance gains remain consistent across diverse models, workloads, and configurations. Code and data are open-sourced.
@@ 1. Introduction
The rise of Large Language Models (LLMs) like GPT-4 (OpenAI et al., 2024), Llama  (Touvron et al., 2023), and Gemini (Team et al., 2024) has revolutionized modern applications, such as chatbots (OpenAI et al., 2024), virtual assistants  (Guan et al., 2023), code generation  (Jiang et al., 2024), and creative writing  (Gómez-Rodríguez and Williams, 2023). This places immense pressure on LLM inference serving systems, which must meet stringent latency requirements for a seamless user experience and to maximize throughput to handle growing demand (Shen and Sen, 2025; Cheng et al., 2025; Kwon et al., 2023; Yu et al., 2022). To this end, in recent years, we have observed significant progress in the development of LLM serving systems. Several key techniques, such as Continuous Batching (Yu et al., 2022), Paged Attention (Kwon et al., 2023), Chunked Prefill (Agrawal et al., 2023), and FlashAttention (Dao, 2023), have been developed with significant improvements in both latency and throughput.
@@
However, the current LLM inference process is often characterized as unpredictable (Sun et al., 2024). For example, the autoregressive nature of LLMs, which generates tokens sequentially based on preceding ones until a stop signal is reached, leads to variable response lengths and decoding steps (Schuurmans et al., 2024). Paged Attention (Kwon et al., 2023), which dynamically allocates memory resources and allows for request preemption, further contributes to the dynamic nature of runtime memory consumption. Furthermore, latency of each decoding step also exhibits high variance, due to dynamic batch size.
@@
The above uncertainties challenge both scheduling LLM requests and scaling model instances, since widely used runtime metrics such as latency, throughput and memory, no longer accurately represent end-to-end execution loads in the case of LLM serving. Further, scheduled requests with unexpectedly long responses can further increase memory load on overloaded hosts and block subsequent requests from being launched (Sun et al., 2024). Most current multi-instance serving frameworks in production, such as (nvi, [n. d.]; Aminabadi et al., 2022), employ scheduling heuristics like round-robin for request dispatching which offer no guarantee on scheduling performance. Alternative solutions, such as Llumnix (Sun et al., 2024), schedule tasks while aligning with dynamic re-balancing through live migration, demonstrating improvements in overall cluster serving performance. However, such solutions require transferring requests’ KV cache (Kang et al., 2024) across instances with extra cost of network bandwidth and memory, which may not be preferable in a resource-constrained cluster with high serving pressure.
@@
Recent studies indicate that the uncertainty of inference behavior can be mitigated with the help of purposely built assistant models. For example, the length of responses and duration of execution can be accurately predicted with a pretrained regression model (Zheng et al., 2023b), and via the inference simulation framework (Agrawal et al., 2024a). This presents an unexplored opportunity to improve load balancing in inference systems by exploiting such approaches. For instance, consider the request to “explain the theory of relativity”, which involves short prompts but generates lengthy responses. Predicting a request’s length allows the scheduler to proactively route it to less-loaded devices, which would improve cluster balance.
@@
In this paper, we present Block, a novel decentralized scheduler for LLM inference clusters. The key novelty of Block relies in leveraging queries’ context and utilizing static properties of LLM serving backends, such as hardware capability and serving framework batching strategy, to predictively schedule inference requests to serving instances.
@@
Block operates in the following way. A lightweight LLM-based regression model is first applied to estimate the length of response based on the request contexts. Then, a simulation framework is applied to predict key target metrics for each request. Finally, the Block task scheduler dispatches requests based on these simulation-based predictions. Following the above steps, Block can estimate real serving load and metrics for each request during scheduling, thereby improving overall performance by maximizing both throughput and resource utilization while minimizing both latency and preemptions. Additionally, our work identifies that such simulation-based prediction provides an efficient approach for auto-provisioning.
@@
We evaluate Block against widely-used heuristic dispatchers on a 12 GPUs cluster with real-world datasets and LLMs. Results demonstrate Block’s improved performance: it increases serving capacity by up to 16.7% and boosts throughput by up to 4.4% against the Llumnix dispatcher. Critically, Block reduces average request latency by 19.9-45.8% and P99 tail latency by 12.6-49.5% compared to baselines. Block’s performance gains are more pronounced when looking at the Time-To-First-Token (TTFT) where average and P99 TTFT are reduced by 88.1-97.0% and 78.6-94.5% respectively. Further, Block improves resource utilization and achieves a 20.1% reduction in P99 latency when using prediction for auto-provisioning.
@@
In summary, this paper makes the following contributions. First, we thoroughly motivate our approach to using predictive scheduling for LLM serving with the integration of response length estimation and inference simulation techniques in §2. Second, we design (§4) and implement(§5) Block, a distributed scheduling framework with predictive scheduling. Third, we conduct a comprehensive evaluation on a 12 GPUs cluster with real-world datasets/models to demonstrate the superior effectiveness of Block on load balancing and auto-provisioning, as presented in §6. Finally, Block paves the way for a new class of predictive LLM schedulers for online serving. Block’s code and dataset are open-sourced at https://github.com/AKafakA/Block.
@@ 2. Background
Modern large language models (LLMs) are mainly built on transformer architectures (Vaswani et al., 2017) to process input sequences. LLM queries typically consist of strings of variable lengths known as prompts or contexts. Queries are first tokenized and converted into a sequence of embeddings, which are then processed by transformer blocks (Vaswani et al., 2017) by applying attention mechanisms in conjunction with multilayer perceptrons to project them into a specified space. The procedure of processing prompts is known as encoding.
@@
Due to autoregressive generation, the decoding phase cannot be accelerated through parallelization and requires caching all tokens until completed. The only exception occurs at the first decoded token, as it only uses the prompt as input without depending on a prior stage. This phase, which involves computing the KV cache for the context sequence, is referred to as the prefill phase, in contrast to the subsequent sequential decoding steps known as the decode phase. To this end, prefill is considered as compute-bound while decoding is regarded as memory-bandwidth-bound (Kamath et al., 2024).
Continuous Batching and Paged Attention
@@
Figure 1. Paged Attention and Preemption
A key challenge in Continuous Batching is memory allocation. Since response length is unknown at runtime, memory must be pre-allocated for the maximum possible sequence length and leads to memory fragmentation and wastage, especially for requests generating shorter responses. This inefficiency on memory utilization necessitates a more dynamic memory management solution. To address this challenge, Paged Attention (Kwon et al., 2023) was proposed and open-sourced through the vLLM serving framework. In the case of Paged Attention, instead of reserving maximum memory or risking an OutOfMemory error, vLLM divides GPU memory into fixed-size memory blocks and maintains a page table to track the locations of noncontinuous physical memory blocks for KV cache associated with each request. As illustrated in Figure 1, preemption occurs when memory is insufficient for the next decoding step. The newest request in the batch is preempted, returned to the head of the waiting queue, and its memory blocks are released for the remaining requests. Once memory becomes available, typically from completed requests, the preempted request resumes, requiring its KV cache to be recomputed. This dynamic memory allocation significantly improves the LLM serving instance’s capacity and throughput, making vLLM a leading framework in research and industry applications.
@@ 3. Related Work and Motivation
This inherent unpredictability in resource requirements and execution behavior makes dispatching requests across LLM serving instances more challenging than generally dynamic task scheduling (Savvas and Kechadi, 2004). For example, a standard scheduler like Kubernetes might filter compute nodes based on initially available resources and then selects the node with the lowest current utilization. However, such static filtering and simple utilization metrics are often inadequate for LLM inference workloads because memory demands of requests are dynamic, varying with the decoding length. Additionally, the actual performance when executing a request can still be impacted by interference from other colocated requests (Sun et al., 2024). Furthermore, unexpected preemption can also occur leading to serving request reset and successors blocked.
@@
To ensure a stable application and a good user experience in serving clusters with variable workloads, effective resource management must go beyond scheduling to include auto-provision to handle peak loads and guarantees performance under established SLOs. Although auto-provision is highly desirable and comes with well-known challenges often faced in other areas with similar scheduling characteristics, yet it is often overlooked in LLM serving systems design. Asynchronous cold starts in serverless computing (Liu et al., 2023) exemplify the issue. While a new instance is being provisioned and initialized, incoming requests continue to be routed to existing instances, exacerbating tail latency. The effect is even more pronounced for LLM serving. In addition to new arrivals, ongoing inference requests also continue generating tokens and consuming additional memory on already overloaded instances. This prolongs latency and increases load, even as the newly provisioned instances remain underutilized. Finally, the resulting load imbalance wastes resources and degrades serving performance.
Llumnix (Sun et al., 2024) can mitigate such imbalances through dynamic re-balancing. Initially, it dispatches requests using a heuristic scheduler. Subsequently, Llumnix performs continuous dynamic load rebalancing across instances by migrating active requests along with their KV caches. While dynamic rebalancing can mitigate runtime load imbalances, it requires significant GPU memory and inter-GPU network bandwidth to transfer associated token caches, impacting model parallelism performance. This resource contention can even intensify in either: 1) when auto-provisioning is triggered due to cold start issues, to re-balance load between heavily utilized existing hosts and newly provisioned ones; or 2) when other network-heavy features are applied, such as P-D disaggregation and tensor/pipeline parallelism (Shoeybi et al., 2020; Huang et al., 2019).
@@
With the rapid development and expansion of new LLM models, devices, and use cases, setting up an appropriate LLM serving cluster has become a critical yet challenging task for developers, particularly due to the high trial-and-error costs associated with GPU pricing. To address this, Vidur, the first LLM cluster simulation framework, has been proposed (Agrawal et al., 2024a). It aims to reduce potential hardware costs when searching for the optimal cluster configuration based on a given model, trace, and Service Level Objective (SLO) requirements. The insight behind Vidur is that since the local scheduler and batching logic are deterministic, and response lengths are typically available in the replay trace used to evaluate cluster performance, it is feasible to simulate the entire replay process if the execution time for each batch can be inferred. Additionally, by profiling low-level operators like attention and linear projection for a specific GPU, Vidur trains linear models to interpolate execution times for various batches. It finally achieves less than 9% error for key metrics such as throughput and latencies across a diverse set of models and GPU types. Besides, SimAI (Wang et al., 2025) also provides modeling and simulation for the entire LLM training process with average 98% accuracy.
@@
While some methods predict inference performance assuming known response lengths, alternative approaches also predict the initially unknown lengths to enhance inference itself. Recognizing that the variable length of responses is a key challenge, these techniques leverage prediction to improve local scheduler efficiency. For instance, Sequence Scheduling (Zheng et al., 2023b) proposes instruction tuning a smaller LLM specifically to predict the response length of the main serving LLM. This prediction allows their local scheduler to be optimized by grouping requests with similar expected lengths into the same batch. Such grouping minimizes computational waste and improves single instances’ throughput by 85%.
Similarly, LightLLM (Gong et al., 2025) predicts request output lengths using historical distributions. Based on these predictions, it infers the peak memory requirement for running batches, enabling the local scheduler to proactively avoid preemption. In the context of global scheduling, length estimation has also been utilized in recent works, primarily as a filtering mechanism prior to scheduling tasks. For instance, TetriServe (Hu et al., 2024) deploys a lightweight length-estimation model to filter out decoding candidate instances without sufficient GPU memory to accommodate the estimated number of tokens and then applies the Power-of-Two method (Richa et al., 2000) to select decoding instances with fewer pending requests. DynamoLLM (Stojkovic et al., 2024) classifies requests into three distinct pools based on their estimated lengths, scheduling them separately with associated host pools. To summarize, our insight is that by combining simulation techniques with length prediction, the unpredictability can be largely eliminated, which in turn creates an opportunity to improve multi-instance LLM serving performance.
@@ 4. Block System Design
Block, as shown in Figure 4, comprises four services: query length tagger, global scheduler, predictor, and inference framework backend, to handle request distribution and results collection. The query length tagger service is the entry point of the system and is designed to predict and label the anticipated response lengths to requests. LLM Requests are executed on the GPUs running under an inference framework such as (Kwon et al., 2023; Zheng et al., 2024; Gong et al., 2025). The global scheduler is tasked with making scheduling and auto-provisioning decisions and directing requests to appropriate model instances to balance the load. When a new query is received, the global scheduler first disseminates prediction requests to the predictor services in order to collect instance statuses/predictions for scheduling. Then, it dispatches requests to the selected model instances. Further, the predictor service performs as a sidecar to predict the target metrics. The Block framework is designed to be agnostic to models, hardware, inference frameworks, and scheduling strategies.